# Bright and dark breathers on an elliptic wave in the defocusing mKdV equation

### v1

@@ 1.4. Methods and organization of the paper
which relates solutions $u(x,t)$ of the mKdV equation (1.1) to solutions $v(x,t)$ of the KdV equation $v_{t}+6vv_{x}+v_{xxx}=0$. A similar transformation maps potentials of the spectral problem (1.3) to potentials of the stationary Schrödinger equation. By using Weierstrass’ elliptic functions, we show that the Schrödinger equation for the general elliptic wave (1.8) becomes the Lamé equation with a single-gap elliptic potential, solutions of which are well-known, see [28, p.395] and [22, Section 3]. . This yields the representation of the eigenfunctions of the spectral problem (1.3) as a linear combination of two solutions of the Schrödinger equation. The coefficients of this linear combination are uniquely computed by using expansions of elliptic functions near poles in the complex plane. This is achieved in Section 2 with the complex parameter $\gamma\in[0,K]\times[0,iK^{\prime}]$ to represent the eigenfunctions in the form (1.16) and (1.17).
@@
Breathers in the form (1.26) are studied in Section 4, where we justify the conditions (1.27) and (1.28) as well as the choice (1.30) for the breather solution to be real-valued and bounded. For the positive half-gap $(0,\zeta_{3})$, where bright breathers of Figures 3 and 4 are constructed, we only need to rewrite the eigenfunctions of Theorem 1 for $\gamma=iK^{\prime}+\delta$ with $\delta\in(2\alpha,K)$ and choose the norming constants conveniently. However, for the positive gap $(\zeta_{2},\zeta_{1})$, where dark breathers of Figures 5 and 6 are constructed, we use $\gamma\in(0,K)$ but change $x$ to $iK^{\prime}+x$. This transformation maps the bounded profile $\phi$ of the elliptic wave to the singular profile but also maps the singular profile of the breather solution $u=u(x,t)$ into a bounded profile.
Finally, in Section 5, we discuss the factorization formula for elliptic eigenfunctions of the Lax system (1.2) associated with the elliptic wave in the factorized forms (1.12) or (1.13). We show that the poles and zeros of the elliptic eigenfunctions are defined by using the branch pole singularities, which cannot be unfolded with the use of elliptic functions.
@@ 1.5. Notations for elliptic functions
where $q:=e^{-\frac{\pi K^{\prime}}{K}}$ with $K$ being the complete elliptic integral for elliptic modulus $k$ and $K^{\prime}$ being the same for $k^{\prime}=\sqrt{1-k^{2}}$. For notational convenience, we use
@@
We derive the explicit representations (1.16) and (1.17) of elliptic eigenfunctions in Theorem 1. To do so, we first give the relation between the elliptic wave (1.8) and Weierstrass’ elliptic functions (Section 2.1), which is used in the derivation. Then, we proceed with the Miura transformation (1.32) to derive the explicit solutions of the Lamé equation with a single-gap elliptic potential (Section 2.2). Finally, we obtain the unique expressions for the linear superpositions of explicit solutions of the Lamé equation from expansion of elliptic functions near poles in the complex plane (Section 2.3).
@@ 2.1. Relation to Weierstrass’ elliptic function
$\left\{\begin{array}[]{l}\displaystyle e_{1}=\frac{1}{3}(\zeta_{1}^{2}+\zeta_{2}^{2}-2\zeta_{3}^{2}),\\ \displaystyle e_{2}=\frac{1}{3}(\zeta_{1}^{2}-2\zeta_{2}^{2}+\zeta_{3}^{2}),\\ \displaystyle e_{3}=\frac{1}{3}(-2\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2}),\end{array}\right.\quad\Rightarrow\quad\left\{\begin{array}[]{l}e_{1}-e_{2}=\zeta_{2}^{2}-\zeta_{3}^{2},\\ e_{1}-e_{3}=\zeta_{1}^{2}-\zeta_{3}^{2},\\ e_{2}-e_{3}=\zeta_{1}^{2}-\zeta_{2}^{2}.\end{array}\right$. (2.2)
@@ 2.2. Miura transformation
$\left(\begin{matrix}-\partial_{x}^{2}+\phi^{2}&\phi^{\prime}\\ \phi^{\prime}&-\partial_{x}^{2}+\phi^{2}\end{matrix}\right)\psi=\zeta^{2}\psi,$ (2.9)
where $\psi=(p,q)^{T}$ is the eigenfunction and $\zeta\in\mathbb{R}$ is the spectral parameter. By folding (2.9) in variables $\psi_{\pm}=p\pm q$, we get two uncoupled Schrödinger equations
@@
Recall from [28, p.395] that the general solution of the Lamé equation
@@
is given by
@@
where $c_{1},c_{2}\in\mathbb{C}$ are arbitrary coefficients and $\gamma\in[-K,K]\times[-iK^{\prime},iK^{\prime}]$ is defined from the spectral parameter $\eta\in\mathbb{R}$ by the dispersion relation $\eta=k^{2}+\operatorname{dn}^{2}(\gamma)$.
Since $z=\nu x\pm\alpha$, we obtain by comparing (2.14) with (2.15) that
where $c_{1}^{\pm},c_{2}^{\pm}\in\mathbb{C}$ are arbitrary coefficients, the factor $2$ is used for convenience, and $\gamma\in[0,K]\times[0,iK^{\prime}]$ is defined from the spectral parameter $\zeta\in\mathbb{R}$ by the dispersion relation
$\zeta^{2}=\zeta_{1}^{2}-(\zeta_{1}^{2}-\zeta^{2})\operatorname{sn}^{2}(\gamma)=\zeta_{2}^{2}+(\zeta_{1}^{2}-\zeta^{2})\operatorname{cn}^{2}(\gamma)=\zeta_{3}^{2}+(\zeta_{1}^{2}-\zeta_{3}^{2})\operatorname{dn}^{2}(\gamma),$ (2.17)
where the fundamental elliptic relations (1.35) have been used. Lemma 1 below clarifies the admissible values of $\gamma\in[0,K]\times[0,iK^{\prime}]$ from the dispersion relation (2.17).
@@ 2.3. Coefficients of the linear superposition
Since $\psi_{\pm}=p\pm q$, we get two linearly independent solutions $\psi=(p_{1},q_{1})$ and $\psi=(p_{2},q_{2})$ of the spectral problem (1.3) from solutions (2.16). The first solution is
@@
Since there exist only two linearly-independent solutions of the spectral problem (1.3), the parameters $(c_{1}^{+},c_{1}^{-})$ and $(c_{2}^{+},c_{2}^{-})$ cannot be arbitrary but must be constant proportional to scalar arbitrary parameters.
@@
The expression (2.18) yields (1.20) for $\gamma=2\alpha+iK^{\prime}$ if $c_{1}^{+}=0$ since
@@
The expression (2.19) yields (1.21) for $\gamma=2\alpha+iK^{\prime}$ if $c_{2}^{-}=0$ since
@@
To find the relations between $c_{1}^{+}$ and $c_{1}^{-}$ in the first solution (2.18) for other values of $\zeta$, we compute the pole contributions of the elliptic solutions of the spectral problem (1.3) at
@@
$\frac{H(\nu x+\gamma-\alpha)}{\Theta(\nu x-\alpha)}e^{-\nu xZ(\gamma)}$ $=e^{(\alpha+iK^{\prime})Z(\gamma)}\frac{H(\gamma-2\alpha-iK^{\prime})}{\Theta(-2\alpha-iK^{\prime})}+\mathcal{O}\left(x-\frac{v}{2}\right),$
@@
we obtain similarly another linear equation
@@
$=\zeta_{1}^{2}-(\zeta_{1}^{2}-\zeta_{3}^{2})k^{2}\operatorname{sn}^{2}(\gamma)$
$=\zeta_{3}^{2}+(\zeta_{1}^{2}-\zeta_{3}^{2})\operatorname{dn}^{2}(\gamma),$
@@
Recall from Example 2 that $\gamma=2\alpha+iK^{\prime}$ for $\zeta=0$, which yields $c_{1}^{+}=0$ since $\Theta(2\alpha-\gamma)=0$ and $\Theta(2\alpha+\gamma)\neq 0$. Therefore, we use (2.26) of the linear system (2.25) and (2.26) to define up to the constant multiplication factor:
$c_{1}^{+}=-i\zeta,\quad c_{1}^{-}=\frac{\nu\sqrt{k}\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)H(2\alpha)}=\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)},$
where we have used again that $\nu=\zeta_{1}\operatorname{sn}(2\alpha)$. This gives the first solution (2.18) in the final form (1.16).
Similarly, we expand the second solution (2.19) as $x\to\pm\frac{v}{2}$. Here, we use the fact that (2.19) is obtained from (2.18) by replacing $\gamma$ to $\gamma$. Since the poles $\pm\frac{v}{2}$ are independent of $\gamma$, we obtain the linear system of equations for $(c_{2}^{+},c_{2}^{-})$ by replacing $\gamma$ to $\gamma$ in (2.25) and (2.26):
$i\zeta\frac{c_{2}^{+}\Theta(\gamma)}{\nu\sqrt{k}\Theta(0)}-c_{2}^{-}\frac{\Theta(2\alpha+\gamma)}{H(2\alpha)}=0.$ (2.28)
@@
Again, it follows from Example 2 for $\zeta=0$ that $c_{2}^{-}=0$ since $\Theta(2\alpha-\gamma)=0$ and $\Theta(2\alpha+\gamma)\neq 0$. Therefore, we use (2.28) of the linear system (2.28) and (2.29) to define up to the constant multiplication factor:
$c_{2}^{+}=\frac{\nu\sqrt{k}\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)H(2\alpha)}=\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)},\quad c_{2}^{-}=i\zeta.$
This gives the second solution (2.19) in the final form (1.17).
@@ 3. Time evolution of the eigenfunctions
We derive the time evolution (1.15) of the elliptic eigenfunctions in Theorem 1. To do so, we first inspect the relationship between the spectral parameter $\zeta\in\mathbb{R}$ and the shift parameter $\gamma\in[0,K]\times[0,iK^{\prime}]$ given by the dispersion relation (1.14) (Section 3.1). Then, we separate variables in the Lax system (1.2) and obtain the characteristic polynomial for the traveling waves (Section 3.2), which is also solved by the elliptic functions. Finally, we derive the unique expression (1.18) for $\mu$ in terms of the shift parameter $\gamma$ (Section 3.3).
@@ 3.1. Dispersion relation for eigenfunctions
For $\zeta\in[\zeta_{2},\zeta_{1}]$, we use $\gamma\in[0,K]$. It follows from
@@
Since $\operatorname{cn}^{2}(K)=0$, the image of $\gamma=K+iK^{\prime}$ is $\zeta^{2}=\zeta_{3}^{2}$. By (1.23), we have
$\operatorname{cn}(2\alpha)=\frac{\zeta_{3}}{\zeta_{1}},\quad\operatorname{sn}(2\alpha)=\frac{\nu}{\zeta_{1}},$
and since $\nu^{2}=\zeta_{1}^{2}-\zeta_{3}^{2}$, the image of $\gamma=2\alpha+iK^{\prime}$ is $\zeta=0$. For $\delta\in(0,K)$, we have $\zeta\in(0,\zeta_{3})$. ∎

### v2

@@ 1.4. Methods and organization of the paper
which relates solutions $u(x,t)$ of the mKdV equation (1.1) to solutions $v(x,t)$ of the KdV equation $v_{t}+6vv_{x}+v_{xxx}=0$. The Miura transformation can be used to map the spectral problem (1.3) to the stationary Schrödinger equation. By using Weierstrass’ elliptic functions, we show that the Schrödinger equation for the general elliptic wave (1.8) becomes Lamé equation with a single-gap elliptic potential, solutions of which are well-known, see [31, p.395] and [23, Section 3]. This yields the representation of the eigenfunctions of the spectral problem (1.3) as a linear combination of two solutions of the Schrödinger equation. The coefficients of this linear combination are uniquely computed by using expansions of elliptic functions near poles in the complex plane. This is achieved in Section 2 with the complex parameter $\gamma\in[0,K]\times[0,iK^{\prime}]$ to represent the eigenfunctions in the form (1.16) and (1.17).
@@
Breathers in the form (1.26) are studied in Section 4, where we justify the conditions (1.27), (1.28), and (1.30) for the breather solution to be real-valued and bounded. For the positive half-gap $(0,\zeta_{3})$, where bright breathers of Figures 3 and 4 are constructed, we only need to rewrite the eigenfunctions of Theorem 1 for $\gamma=iK^{\prime}+\delta$ with $\delta\in(2\alpha,K)$ and choose the norming constants appropriately. However, for the positive gap $(\zeta_{2},\zeta_{1})$, where dark breathers of Figures 5 and 6 are constructed, we use $\gamma\in(0,K)$ but change $x$ to $iK^{\prime}+x$. This transformation maps the bounded profile $\phi$ of the elliptic wave in (1.8) to the singular profile but also maps the singular profile of the breather solution $u=u(x,t)$ in (1.26) into a bounded profile.
Finally, in Section 5, we discuss the factorized form for elliptic eigenfunctions of the Lax system (1.2) associated with the elliptic wave. We show that the poles and zeros of the elliptic eigenfunctions depend on the spectral parameter via the branch pole singularities, which cannot be unfolded with the use of elliptic functions.
@@ 1.5. Notations for elliptic functions
where $q:=e^{-\frac{\pi K^{\prime}}{K}}$ with $K$ and $K^{\prime}$ being the complete elliptic integrals for elliptic modulus $k$ and $k^{\prime}=\sqrt{1-k^{2}}$, respectively. For notational convenience, we use
@@
We derive here the explicit representations (1.16) and (1.17) of elliptic eigenfunctions in Theorem 1. To do so, we first give the relation between the elliptic wave (1.8) and Weierstrass’ elliptic functions (Section 2.1), which is used in the derivation. Then, we proceed with the Miura transformation (1.32) to derive the explicit solutions of the Lamé equation with a single-gap elliptic potential (Section 2.2). Finally, we obtain the unique expressions for the linear superpositions of explicit solutions of the Lamé equation from expansion of elliptic functions near poles in the complex plane (Section 2.3).
@@ 2.1. Relation to Weierstrass’ elliptic function
$\left\{\begin{array}[]{l}\displaystyle e_{1}=\frac{1}{3}(\zeta_{1}^{2}+\zeta_{2}^{2}-2\zeta_{3}^{2}),\vskip 7.11317pt\\ \displaystyle e_{2}=\frac{1}{3}(\zeta_{1}^{2}-2\zeta_{2}^{2}+\zeta_{3}^{2}),\vskip 7.11317pt\\ \displaystyle e_{3}=\frac{1}{3}(-2\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2}),\end{array}\right.\quad\Rightarrow\quad\left\{\begin{array}[]{l}e_{1}-e_{2}=\zeta_{2}^{2}-\zeta_{3}^{2},\vskip 2.84544pt\\ e_{1}-e_{3}=\zeta_{1}^{2}-\zeta_{3}^{2},\vskip 2.84544pt\\ e_{2}-e_{3}=\zeta_{1}^{2}-\zeta_{2}^{2}.\end{array}\right$. (2.2)
@@ 2.2. Miura transformation
$\left(\begin{matrix}-\partial_{x}^{2}+\phi^{2}&\phi^{\prime}\\ \phi^{\prime}&-\partial_{x}^{2}+\phi^{2}\end{matrix}\right)\varphi=\zeta^{2}\varphi,$ (2.9)
where $\varphi=(p,q)^{T}$ is the eigenfunction and $\zeta\in\mathbb{R}$ is the spectral parameter. By folding (2.9) in variables $\psi_{\pm}=p\pm q$, we get two uncoupled Schrödinger equations
@@
Explicit solutions of the Schrödinger equations (2.10) are described in the following lemma.
Lemma 1.
Let $\gamma\in[0,K]\times[0,iK^{\prime}]$ be defined from the spectral parameter $\zeta\in\mathbb{R}$ by the dispersion relation
$\zeta^{2}=\zeta_{1}^{2}-(\zeta_{1}^{2}-\zeta_{2}^{2})\operatorname{sn}^{2}(\gamma)=\zeta_{2}^{2}+(\zeta_{1}^{2}-\zeta_{2}^{2})\operatorname{cn}^{2}(\gamma)=\zeta_{3}^{2}+(\zeta_{1}^{2}-\zeta_{3}^{2})\operatorname{dn}^{2}(\gamma).$ (2.11)
Assume that $\gamma\neq\{0,K,iK^{\prime},K+iK^{\prime}\}$. General solutions to the Schrödinger equations (2.10) are given by
where $c_{1}^{\pm},c_{2}^{\pm}\in\mathbb{C}$ are arbitrary coefficients and the factor $2$ is used for convenience.
Proof.
@@
Recall from [31, p.395] that the Lamé equation
@@
admits for $\eta\neq\{1,k^{2},1+k^{2},\infty\}$ a general solution in the form:
@@
where $c_{1},c_{2}\in\mathbb{C}$ are arbitrary coefficients and $\gamma\in[0,K]\times[0,iK^{\prime}]$ is defined from the spectral parameter $\eta\in\mathbb{R}$ by the dispersion relation $\eta=k^{2}+\operatorname{dn}^{2}(\gamma)$. If $\eta\neq\{1,k^{2},1+k^{2},\infty\}$, then $\gamma\neq\{0,K,iK^{\prime},K+iK^{\prime}\}$. Since $z=\nu x\pm\alpha$, we obtain (2.11) and (2.12) by comparing (2.16) with (2.17) and using (1.9) for
$\eta=1+\frac{\zeta^{2}-\zeta_{2}^{2}}{\zeta_{1}^{2}-\zeta_{3}^{2}}.$
The three relations in (2.11) are equivalent due to the fundamental relations for elliptic functions (1.35). ∎
@@ 2.3. Coefficients of the linear superposition
Since $\psi_{\pm}=p\pm q$, we get two linearly independent solutions $\varphi=(p_{1},q_{1})$ and $\varphi=(p_{2},q_{2})$ of the spectral problem (1.3) from solutions (2.12) by using the inverse transformation
$p=\frac{\psi_{+}+\psi_{-}}{2},\quad q=\frac{\psi_{+}-\psi_{-}}{2}.$
Separating the two solutions in (2.12) according to the two exponential factors, we obtain the two linearly independent solutions in the form:
@@
Since there exist only two linearly-independent solutions of the spectral problem (1.3), each set $(c_{1}^{+},c_{1}^{-})$ and $(c_{2}^{+},c_{2}^{-})$ cannot be arbitrary but must be linearly dependent.
@@
The expression (2.18) with $\gamma=2\alpha+iK^{\prime}$ yields (1.20) if $c_{1}^{+}=0$ since
@@
The expression (2.19) with $\gamma=2\alpha+iK^{\prime}$ yields (1.21) if $c_{2}^{-}=0$ since
@@
The following lemma describes the linear relations in each set $(c_{1}^{+},c_{1}^{-})$ and $(c_{2}^{+},c_{2}^{-})$.
Lemma 2.
If the representations (2.18) and (2.19) are solutions of the spectral problem (1.3), then their coefficients are uniquely defined up to the constant multiplication factor as
$c_{1}^{+}=-i\zeta,\quad c_{1}^{-}=\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)}$ (2.20)
and
$c_{2}^{+}=\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)},\quad c_{2}^{-}=i\zeta.$ (2.21)
Proof.
To obtain (2.20), we compute the pole contributions of the elliptic solutions of the spectral problem (1.3) at
@@
$\frac{H(\nu x+\gamma-\alpha)}{\Theta(\nu x-\alpha)}e^{-\nu xZ(\gamma)}$ $=e^{(\alpha+iK^{\prime})Z(\gamma)}\frac{H(\gamma-2\alpha-iK^{\prime})}{\Theta(-2\alpha-iK^{\prime})}+\mathcal{O}\left(x-\frac{v}{2}\right)$
@@
we obtain another linear equation
@@
$=\zeta_{1}^{2}-(\zeta_{1}^{2}-\zeta_{2}^{2})\operatorname{sn}^{2}(\gamma),$
@@
Recall from Example 2 that $\gamma=2\alpha+iK^{\prime}$ for $\zeta=0$, which yields $c_{1}^{+}=0$ since $\Theta(2\alpha-\gamma)=0$ and $\Theta(2\alpha+\gamma)\neq 0$. Therefore, we use (2.28) of the linear system (2.27) and (2.28), set $c_{1}^{+}=-i\zeta$ and obtain
$c_{1}^{-}=\frac{\nu\sqrt{k}\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)H(2\alpha)}=\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)},$
where we have used again that $\nu=\zeta_{1}\operatorname{sn}(2\alpha)$. This yields (2.20).
To obtain (2.21) with a similar method, we expand the second solution (2.19) as $x\to\pm\frac{v}{2}$. Here, we use the fact that (2.19) is obtained from (2.18) by replacing $\gamma$ to $\gamma$. Since the poles $\pm\frac{v}{2}$ are independent of $\gamma$, we obtain the linear system of equations for $(c_{2}^{+},c_{2}^{-})$ by replacing $\gamma$ to $\gamma$ in (2.27) and (2.28):
$i\zeta\frac{c_{2}^{+}\Theta(\gamma)}{\nu\sqrt{k}\Theta(0)}-c_{2}^{-}\frac{\Theta(2\alpha+\gamma)}{H(2\alpha)}=0$ (2.30)
@@
Again, it follows from Example 2 for $\zeta=0$ that $c_{2}^{-}=0$ since $\Theta(2\alpha-\gamma)=0$ and $\Theta(2\alpha+\gamma)\neq 0$. Therefore, we use (2.30) of the linear system (2.30) and (2.31), set $c_{2}^{-}=i\zeta$ and obtain
$c_{2}^{+}=\frac{\nu\sqrt{k}\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)H(2\alpha)}=\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)}.$
This yields (2.21). ∎
The first solution (2.18) with (2.20) yields the final form (1.16) in Theorem 1, whereas the second solution (2.19) with (2.21) yields the final form (1.17).
@@ 3. Time evolution of the eigenfunctions
Having proven the representation (1.16) and (1.17) in Section 2, we derive here the time evolution (1.15) of the elliptic eigenfunctions stated in Theorem 1. To do so, we first inspect the relationship between the spectral parameter $\zeta\in\mathbb{R}$ and the shift parameter $\gamma\in[0,K]\times[0,iK^{\prime}]$ given by the dispersion relation (1.14) (Section 3.1). Then, we separate variables in the Lax system (1.2) and obtain the characteristic polynomial for the traveling waves (Section 3.2), which is also solved by the elliptic functions. Finally, we derive the unique expression (1.18) for $\mu$ in terms of the shift parameter $\gamma$ (Section 3.3).
@@ 3.1. Dispersion relation for eigenfunctions
For $\zeta\in[\zeta_{2},\zeta_{1}]$, we use $\gamma\in[0,K]$ in (1.14), rewritten again as
@@
Since $\operatorname{cn}^{2}(K)=0$, the image of $\gamma=K+iK^{\prime}$ is $\zeta^{2}=\zeta_{3}^{2}$. By (1.23) and $\nu^{2}=\zeta_{1}^{2}-\zeta_{3}^{2}$, the image of $\gamma=2\alpha+iK^{\prime}$ is $\zeta=0$. For $\delta\in(0,K)$, we have $\zeta\in(0,\zeta_{3})$. ∎